# Video Super-Resolution: All You Need is a Video Diffusion Model

### v2

@@ Abstract
We present a generic video super-resolution algorithm in this paper, based on the Diffusion Posterior Sampling framework with an unconditional video generation model in latent space. The video generation model, a diffusion transformer, functions as a space-time model. We argue that a powerful model, which learns the physics of the real world, can easily handle various kinds of motion patterns as prior knowledge, thus eliminating the need for explicit estimation of optical flows or motion parameters for pixel alignment. Furthermore, a single instance of the proposed video diffusion transformer model can adapt to different sampling conditions without re-training. Empirical results on synthetic and real-world datasets demonstrate that our method has strong capabilities to address video super-resolution challenges.
Keywords Video super-resolution  $\cdot$ diffusion transformer  $\cdot$ video diffusion model
1 Introduction
@@
    By integrating a transformer network as the denoiser in the reverse diffusion process, rather than a UNet, the model achieves superior scalability and enhanced effectiveness in managing complex and dynamic scenarios.
@@
This algorithm is founded on our belief that with unconditional DMs, explicit motion estimation over time is unnecessary, akin to how facial symmetry is naturally maintained in face image generation without explicit interventions [30]. However, it demands that the DMs not only learn the distribution of video contents but also understand their dynamics governed by real-world physics. To validate our approach, we employ a synthetic video dataset to explore the behavioral dynamics of the method.
2 Related work
2.1 Diffusion models
@@
Most diffusion model-based SISR and VSR applications utilize the above conditional networks. In such scenarios, if the sensing model undergoes changes (e.g., slight alterations in the lens filter), then the conditional networks typically require retraining.
2.2 DPS
@@
2.3 Video Diffusion Models (VDMs)
@@
3 Proposed Approach
Figure 1: Overview of the Video Diffusion Model based VSR (VDM-VSR): In the core iteration, the estimated 3D HR video resides in the latent space, represented by green boxes. It is generated and refined by the VDM, which includes several Transformer blocks. The latent video is then decoded and compared with the LR observations through the degradation model, indicated by red boxes. The discrepancies between these observations and the latent video are used to correct and enhance the HR video during the iteration. Upon completion of this iteration, the latent video is decoded back to the conventional HR space.
We introduce VDM-VSR, a novel approach that addresses super-resolution as an inverse problem under the DPS framework. This method utilizes a DiT-based VDM as the denoiser. We posit that as long as the VDM effectively learns the dynamics of the world as represented by the training video dataset, it should inherently manage inter-frame motion estimation. An overview of the proposed algorithm’s architecture is depicted in Fig. 1.
@@ 3 Proposed Approach
4 Experiments
4.1 Synthetic Dataset
@@
Figure 5: PSNR v.s. number of used frames. LR inputs are generated with with blur kernel $\sigma_{h}=0,2,4,6,8,10$.
4.2 BAIR Dataset
To evaluate our method on real-world data, we used the BAIR robot pushing dataset [39], which consists of 90K short video clips recorded by a real camera. Although this setting remains somewhat of a “toy world” (featuring robotic arms in a controlled environment), it introduces more natural lighting, scene textures, and frequent occlusions than Moving MNIST. We trained our model using 130K video clips, each video contains 10 frames. Our approach effectively reconstructed sharp videos under these conditions, often achieving near-perfect restoration of the robot arm’s position and scene details for both 4x and 8x down-sampled scenarios (see Figure 6 and Figure 7).
@@ 4.2 BAIR Dataset
We compared our algorithm to two state-of-the-art video super-resolution methods: IART [20] and FMA-Net [18]. For a fair comparison, we retrained these methods on the BAIR robot pushing dataset following their official implementation. We set the down-sampling factor to 4x and measured PSNR and SSIM against the corresponding ground-truth videos for quantitative evaluation (see Table 1). Both methods under comparison are based on motion estimation modules, thus presenting limitations in effectively recovering complex motion dynamics. Our method’s results highlight that the utilization of inter-frame motion cues improves VSR performance while obviating the requirement for explicit motion estimation.
Table 1: Quantitative Comparison on BAIR dataset.
@@
5 Conclusions
In this paper, we introduce a generic VSR algorithm based on the DPS framework, incorporating an unconditional video diffusion model. Unlike the original DPS approach, its reverse diffusion iteration operates in a latent image space extended with a temporal axis, and the denoiser is powered by a transformer neural network. This transformer is pre-trained through an unconditional video diffusion process, enabling it to learn the physics of the world across both spatial and temporal dimensions. We also refined the degradation formula by integrating per-frame downsampling and frame masking to effectively address video super-resolution challenges.

### v4

# Rethinking Video Super-Resolution: Towards Diffusion-Based Methods without Motion Alignment

@@ Abstract
In this work, we rethink the approach to video super-resolution by introducing a method based on the Diffusion Posterior Sampling framework, combined with an unconditional video diffusion transformer operating in latent space. The video generation model, a diffusion transformer, functions as a space-time model. We argue that a powerful model, which learns the physics of the real world, can easily handle various kinds of motion patterns as prior knowledge, thus eliminating the need for explicit estimation of optical flows or motion parameters for pixel alignment. Furthermore, a single instance of the proposed video diffusion transformer model can adapt to different sampling conditions without re-training. Empirical results on synthetic and real-world datasets illustrate the feasibility of diffusion-based, alignment-free video super-resolution.
Index Terms:
Video Super-Resolution, Diffusion Transformer, Video Diffusion Models
I Introduction
@@
    By integrating a transformer network as the denoiser in the reverse diffusion process, the model achieves superior scalability and enhanced effectiveness in managing complex and dynamic scenarios.
@@
This algorithm is founded on our belief that with unconditional DMs, explicit motion estimation over time is unnecessary, akin to how facial symmetry is naturally maintained in face image generation without explicit interventions [30]. However, it demands that the DMs not only learn the distribution of video contents but also understand their dynamics governed by real-world physics. To validate our approach, we employ synthetic and real-world data to explore the behavioral dynamics of the method.
II Related work
II-A Diffusion models
@@
II-B Diffusion Posterior Sampling (DPS)
@@
II-C Video Diffusion Models (VDMs)
@@
III Proposed Approach
Figure 1: Overview of the Video Diffusion Model based VSR (VDM-VSR): In the core iteration, the estimated 3D HR video resides in the latent space, represented by green boxes. It is generated and refined by the VDM, which includes several Transformer blocks, as shown in the structural diagram on the right. The latent video is then decoded and compared with the LR observations through the degradation model, indicated by red boxes. The discrepancies between these observations and the latent video are used to correct and enhance the HR video during the iteration. Upon completion of this iteration, the latent video is decoded back to the conventional HR space.
We introduce VDM-VSR, an approach that addresses super-resolution as an inverse problem under the DPS framework. This method utilizes a DiT-based VDM as the denoiser. We posit that as long as the VDM effectively learns the dynamics of the world as represented by the training video dataset, it should inherently manage inter-frame motion estimation. An overview of the proposed algorithm’s architecture is depicted in Fig. 1.
@@ III Proposed Approach
where $H(\mathbf{X},\mathbf{h})$ is a simplified formulation of the frame-wise degradation process, which approximates the combined effects of spatial convolution and downsampling applied to each frame $\mathbf{x}_{j}$ in the sequence.
@@
IV Experiments
IV-A Synthetic Dataset
@@
Figure 5: PSNR v.s. number of used frames. LR inputs are generated with blur kernel $\sigma_{h}=0,2,4,6,8,10$.
@@
IV-B BAIR Dataset
To evaluate our method on real-world data, we used the BAIR robot pushing dataset [39], which consists of 90K short video clips recorded by a real camera. Although this setting remains somewhat of a “toy world” (featuring robotic arms in a controlled environment), it introduces more natural lighting, scene textures, and frequent occlusions than Moving MNIST. We trained our model using 130K video clips, each video contains 10 frames. Our approach effectively reconstructed sharp videos under these conditions, often achieving near-perfect restoration of the robot arm’s position and scene details for both 4x and 8x down-sampled scenarios (see Fig. 6 and Fig. 7).
We compared our algorithm to two state-of-the-art video super-resolution methods: IART [20] and FMA-Net [18]. For a fair comparison, we retrained these methods on the BAIR robot pushing dataset following their official implementation. We set the down-sampling factor to 4x and measured PSNR and SSIM against the corresponding ground-truth videos for quantitative evaluation (see Table I). Both methods under comparison are based on motion estimation modules, thus presenting limitations in effectively recovering complex motion dynamics. Our method’s results highlight that the utilization of inter-frame motion cues improves VSR performance while obviating the requirement for explicit motion estimation.
TABLE I: Quantitative Comparison on BAIR dataset.
@@ IV-B BAIR Dataset
V Conclusions
In this paper, we introduce a novel VSR algorithm based on the DPS framework, incorporating an unconditional video diffusion model. Unlike the original DPS approach, its reverse diffusion iteration operates in a latent image space extended with a temporal axis, and the denoiser is powered by a transformer neural network. This transformer is pre-trained through an unconditional video diffusion process, enabling it to learn the physics of the world across both spatial and temporal dimensions. We also refined the degradation formula by integrating per-frame downsampling and frame masking to effectively address video super-resolution challenges.